OpenStudy (christian_10_):

What is the value of the expression. (–4)4

OpenStudy (michele_laino):

hint: \[\Large {\left( { - 4} \right)^4} = \left( { - 4} \right) \times \left( { - 4} \right) \times \left( { - 4} \right) \times \left( { - 4} \right) = ...?\]

OpenStudy (christian_10_):

256

OpenStudy (christian_10_):

Okay! I know what to do now! thanks @Michele_Laino :)